# Consumption Partial Insurance

### v3

@@ 4.2 Estimation
Estimation follows a three-step approach. In the first step, we compute income and consumption growth net of their predictable components, respectively, $\mathbf{X}_{it}$ and $\mathbf{Z}_{it}$ in model notation, by regressing real income and consumption growth on a rich set of time effects and household characteristics. [^20] The controls include year and year-of-birth dummies, dummies for education, race, region, family size and number of children (level and change from last period), employment status (current and past), presence of outside dependents, and presence of income recipients other than the husband and wife (both current and past). The dummies for education, race, and region are interacted with the year dummies. The residuals from these regressions constitute the empirical counterparts to unexplained income and consumption growth, $\Delta y_{it}$ and $\Delta c_{it}$, respectively. In the second step, we estimate the parameters of the income process – i.e., the second and higher moments of permanent and transitory shocks – using income data alone. In the third step, we estimate the transmission parameters of shocks and the moments of taste heterogeneity and consumption measurement error; we do so using income and consumption data conditional on the parameters of the income process from the second stage.
@@
Given the multi-step approach, we conduct inference using the block bootstrap jointly over all stages. We thus account for serial correlation and heteroskedasticity of arbitrary form, as well as for our use of pre-estimated parameters in the later stages of the multi-step approach. To avoid the standard errors being affected by extreme draws, we apply a normal approximation to the inter-quartile range of bootstrap replications. [^22] Instead of estimating standard errors as the standard deviation of the parameters’ point estimates across replications, we compute them as the normalized interquartile range of the replications, i.e., the difference between the $75^{\text{th}}$ and $25^{\text{th}}$ percentiles divided by the interquartile range of the standard normal cdf. This is equivalent to applying a normal approximation to the distribution of bootstrap replications, thereby shielding the standard errors from extreme resampling draws. See Blundell et al. (2016) and Theloudis (2021) for details. Standard errors without the normal approximation are very similar.
@@ 4.3 Empirical moments of income and consumption
Table 2 reports the empirical second, third, and fourth moments of the cross-sectional distributions of unexplained income and consumption growth in the sample.
Income moments. Panel A presents the moments of income growth. The variance is $0.153$, while the first-order autocovariance is $-0.049$ and highly significant. [^24] In contrast to the identifying statements of section 3, the data are biennial so we report moments for income/consumption growth at a biennial frequency. We maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that this corresponds to a difference over two years. See appendix B for details. The income process of section 2.1 implies that income growth is negatively first-order autocorrelated due to mean reversion in the transitory shock, and this is supported by the data. [^25] The second-order autocovariance is an order of magnitude smaller and only marginally significant. This feature and the biennial nature of the data, whereby the second-order autocovariance reflects correlations 4 calendar years apart, motivates our white noise specification for the transitory shock.
@@
Notes: The table presents the second, third, and fourth moments of income and consumption growth in the baseline sample. Skewness and kurtosis correspond to the third and fourth standardized moments respectively. Column 1 reports moments of consumption internally in the PSID; column 2 reports moments of consumption imputed from the CEX. See appendix C.3 for details on the imputation. We maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that, given the biennial nature of the data, this corresponds to a difference over two calendar years. Block bootstrap standard errors are in parentheses.
@@
Guvenen et al. (2021) focus on male earnings while our model and statistics for income growth above revolve around household disposable income. To facilitate the comparison, appendix C.1 reports moments of male earnings growth in our sample. The variance is very similar between the PSID and the U.S. Social Security Administration data, at $0.271$ and $0.323$ respectively. Skewness is $-0.569$ (and highly statistically significant) versus $-1.039$ in the administrative data. While the nature of skewness is similar in both cases, the population data feature substantially more left skewness than the PSID; Guvenen et al. (2021) confirm this. Kurtosis, finally, is identical between the two, at $13.711$ and $13.494$ respectively.
Left skewness and excess kurtosis are thus not just features of the growth rates in disposable household income (table 2), but they are also observed in alternative measures of income, such as male earnings (table C.1, panel A) and household earnings (table C.1, panel B). Together, left skewness and excess kurtosis suggest that income changes for most households are either close to their mean value of zero or far out in the left tail (large income reductions). These results corroborate (and, in fact, are very close numerically with) findings by De Nardi et al. (2019); De Nardi et al. (2021) over an earlier period of the PSID.
Large income reductions do not just reflect unemployment, on which there is an abundant existing literature, but a multitude of events. Issues of reverse causality aside, appendix C.2 documents that households with large income reductions also experience the loss of a second job, worsening health, large drops in hours worked (possibly reflecting both demand and supply factors), and precarious housing. It is unclear how persistent those events are, but our subsequent decomposition of income to moments of permanent and transitory shocks will shed light on this.
Consumption moments. Panel B of table 2 presents the moments of consumption growth; column 1 specifically reports moments from our baseline sample in the PSID. The variance of consumption growth is very similar to that of income growth ($0.126$ versus $0.153$), reflecting either a large pass-through of income shocks or large taste heterogeneity/measurement error. The first-order autocovariance is $-0.044$ and highly significant. In the context of our model, this autocovariance reflects the variance of consumption error $\sigma_{u^{c}_{t}}^{2}$ (see appendix B).
@@
Appendix figure C.1 plots the variance, skewness, and kurtosis of log consumption in the PSID and CEX samples. The patterns for the variance and kurtosis are quite similar in the two surveys over time, though the variance in the PSID is lower and the kurtosis is slightly higher. Skewness, however, is markedly different. Skewness of log consumption is strongly positive in the PSID, while it is mostly negative in the CEX; the patterns also diverge over time. Overall, these statistics suggest that consumption across otherwise similar samples in the PSID and CEX exhibits some notable differences over our sample period.
This comparison is not informative about the moments of consumption growth, which is the relevant variable in our case. The CEX lacks the panel dimension needed for our analysis, so it cannot be used to obtain measures of consumption growth. This is why BPP imputed consumption from the CEX into the PSID: the CEX lacked the necessary panel dimension while the PSID lacked comprehensive consumption data before 1999. We repeat BPP’s imputation over the recent years (our sample period 1999–2019). The imputation consists of the estimation of food demand in the CEX (food expenditure expressed as a function of prices and nondurable consumption). The demand equation is then inverted (nondurable consumption expressed as a function of prices and food) and used to impute nondurables in the PSID based on food expenditure and prices that are available in both surveys. In this way, we obtain an alternative, external measure of consumption in the PSID, and of consumption growth. Appendix C.3 provides details on the imputation.
The moments of imputed consumption growth are in column 2 of table 2. The variance at $0.323$ is two-and-a-half times higher than the variance in the true series (column 1). In fact, the variance of imputed consumption growth is more than double the variance of income growth. The higher-order moments also differ markedly between the imputed and true series. Skewness in imputed consumption growth is $-0.177$ (versus $-0.039$ in the true series), while kurtosis is $17.109$, almost four times higher than kurtosis in the true series. The imputation thus seems to impart substantial measurement/imputation error to consumption.
Since the imputation is not strictly needed over our sample period, we use consumption internally available in the PSID as our baseline measure of consumption. The present analysis also suggests that we may not want to rely on higher-order moments of consumption, given the notable discrepancies in skewness of log consumption between the PSID and the CEX (this is independent of the imputation). As we established in section 3.3, an advantage of our identification approach over other methods is precisely its non-reliance on such higher-order consumption moments.
@@ 5 Results
We first present baseline results imposing stationarity in the distributions of income shocks, taste heterogeneity, consumption error, and a constant pass-through of income shocks over the lifecycle. [^28] Stationarity is not an identifying assumption but helps with precision, especially of higher moments. We subsequently relax stationarity/lifecycle-invariance in section 5.4.
@@ 5.1 Income process
We present estimates of the income process in three specifications. First, we target only second moments of income growth, which enables us to estimate the second moments of income shocks. This is similar to what BPP do over the earlier period 1980–1992. Second, we target second and third moments of income growth, which enables us to estimate also the skewness of income shocks. Finally, we target up to the fourth moments of income growth, which allows us to estimate all the parameters of the income process as presented in section 2.1. The results from each specification appear in table 3, columns 1-3, respectively. The full set of moments targeted in each case is shown in appendix B.4.1.
@@
Table 3: Estimates of the income process, baseline sample
@@
Variance $\sigma_{\zeta}^{2}$ 0.027 0.027 0.030
(0.002) (0.002) (0.002)
Skewness $\gamma_{\zeta}$ central -0.004 -0.004
(0.003) (0.003)
standardized -0.947 -0.831
(0.649) (0.571)
Kurtosis $\kappa_{\zeta}$ central 0.039
standardized 44.393
@@
Variance $\sigma_{v}^{2}$ 0.033 0.033 0.031
(0.002) (0.002) (0.003)
Skewness $\gamma_{v}$ central -0.008 -0.008
(0.003) (0.003)
standardized -1.310 -1.431
(0.510) (0.566)
Kurtosis $\kappa_{v}$ central 0.048
standardized 49.226
@@
Targeting the third moments of income growth (in addition to the second moments) leaves the variances unchanged. We estimate the skewness (third standardized moment) at $-0.947$ and $-1.310$ for permanent and transitory shocks, respectively. Both shocks thus exhibit substantial negative skewness, i.e., their distribution has a much longer left than right tail. This is consistent with Geweke and Keane (2000) who find left skewness of earnings shocks in the early PSID, and Theloudis (2021) who finds left skewness of wage shocks in the recent period. Assuming for now that good and bad shocks transmit similarly into consumption, the long left tail implies that bad shocks are more damaging than good ones because they are further away from the zero mean – they are more extreme. This is true for both permanent and transitory shocks. Skewness of income growth, however, is driven mostly by permanent shocks (fully so under stationarity). Therefore, if targeting the third moments of income growth subsequently changes the degree of consumption insurance in the household, this must reflect the implications of permanent rather than of transitory shocks.
Targeting all moments of income growth up to fourth-order leaves the third central moments unchanged and modifies the variances only by little. We estimate the kurtosis (fourth standardized moment) at $44.393$ and $49.226$ respectively. In both cases, the distributions are highly leptokurtic: they are sharp-pointed around the mean, have relatively little mass in the shoulders and very fat tails. This is consistent with Busch and Ludwig (2021), who use the PSID until 2012 and estimate numerically similar kurtosis coefficients for both shocks. Together with left skewness, large excess kurtosis implies that most households experience either negligible shocks (shocks close to the zero mean) or shocks that are quite bad. Ignoring these stark features of income shocks may lead to erroneous inference of the degree of consumption partial insurance, to which we now turn.
@@ 5.2 Linear consumption function
We present estimates of the linear consumption function, first, targeting only second moments of the joint income-consumption distribution and, subsequently, targeting second as well as higher-order moments of the joint distribution. The full set of moments targeted in each case is shown in appendix B.4.2.
@@
What explains the discrepancy in $\phi^{(1)}$ with BPP? As suggested earlier, there are a few important differences between our replication and the original exercise in BPP: new consumption data and a different time period. To help decipher their role, we impute consumption from the CEX into the PSID in the same way BPP did for the earlier period, and we repeat the estimation; see appendix C.3 for details on the imputation. The results are in column 2 of table 4. The transmission parameter of permanent shocks doubles to $\phi^{(1)}=0.288$, while the transmission parameter of transitory shocks remains statistically zero (though at a larger magnitude). Using imputed consumption data thus inflates the pass-through of permanent shocks, substantially lowering the degree of partial insurance. This is likely the byproduct of measurement/imputation error, evident in the variance $\sigma^{2}_{u_{c}}$ of consumption error that more than triples between the true series (column 1) and the imputed data from the CEX (column 2). By contrast, the variance of taste heterogeneity $\sigma^{2}_{\xi}$ remains similar in both cases, and also very close to BPP. The imputation, therefore, seems to bring in large amounts of measurement error, explaining about 1/3 of the difference between our lower $\phi^{(1)}$ and BPP’s higher estimate. We relegate the details of the imputation and a discussion of a mechanism through which the imputation in particular inflates $\phi^{(1)}$ to appendix C.3.
Table 4: Estimates of the consumption function, baseline sample
@@
Although the imputation inflates $\phi^{(1)}$, the inflated value in column 2 is still substantially lower than BPP. The PSID switched in 1997 from annual to biennial frequency. It is thus possible that at the lower frequency certain higher frequency income shocks are smoothed out or undone by subsequent shocks. To assess this, we would need to re-estimate our model on annual growth rates, which is impossible after 1997. However, we can use the original BPP data over 1980–1992 to estimate the model annually versus biennially. We do this in appendix C.4. While we obtain $\phi^{(1)}=0.64$ using annual growth rates (exactly as BPP), this drops to $\phi^{(1)}=0.4$–0.5 using biennial rates. The drop is sharper when we target time-invariant moments, which was recently also shown in Chatterjee et al. (2021). The biennial nature of the modern data thus explains another 1/3 of the difference between our lower $\phi^{(1)}$ and BPP’s higher estimate. [^31] In both annual and biennial cases, $\phi^{(1)}$ measures the pass-through of the permanent shock at an annual rate. This becomes clear in appendix B that shows the estimating equations with biennial data. The difference in the estimates of $\phi^{(1)}$ between the two cases arises because with biennial data we only measure an ‘aggregate’ pass-through within a given biennial period. Consequently, some higher-frequency shocks will be inevitably smoothed out at the observed lower frequency and their transmission will be muted. We attribute the remaining gap to calendar time effects, e.g., changes in taxes/benefits or other insurance channels between 1980–1992 and 1999–2019 (see Borella et al., 2022, for a discussion of such changes). It is worth noting that we are not the first paper to find a low pass-through of permanent shocks in the recent years; Arellano et al. (2023) use PSID data after 2005 and find a pass-through of about $0.3$ (specification closest to ours). If we zoom into the period after 2005, we estimate $\phi^{(1)}=0.21$ (results not shown for brevity), statistically indistinguishable from Arellano et al. (2023)’s estimate.
@@
Introducing higher-order moments. The estimation of the linear consumption function targeting higher-order moments is our first attempt to uncover how the tails of income risk affect the pass-through of shocks into consumption. In other words, we want to measure the degree of partial insurance going beyond the workhorse income-consumption covariance to account for the non-Gaussian features of modern income dynamics. We introduce the higher-order moments incrementally. [^32] We use the corresponding pre-estimated income parameters from columns 2 and 3 of table 3 respectively.
@@
Column 3 of table 4 presents the results from targeting third- in addition to second-order moments of consumption and income. The partial insurance parameters remain largely unchanged upon inclusion of the third moments. We estimate $\phi^{(1)}=0.160$ (as opposed to $0.152$ with second moments only) and $\psi^{(1)}=-0.012$ (again indistinguishable from zero). We thus again find large partial insurance to permanent shocks (a $10\%$ permanent change in income results in only $1.6\%$ shift in consumption) and full insurance to transitory shocks. Skewness in taste heterogeneity and consumption error is small (much smaller than skewness in income shocks), while their variances remain unchanged from the specification with second moments only. Therefore, left skewness in consumption is driven by negatively skewed income shocks rather than by taste heterogeneity or consumption measurement error.
@@
Column 4 of table 4 presents the results from targeting all income and consumption moments up to fourth-order. The earlier findings hold up. We estimate $\phi^{(1)}=0.112$ (versus $0.152$ and $0.160$ in the lower-order cases) and $\psi^{(1)}=0.007$ (indistinguishable from zero). A $10\%$ permanent change in income thus results in $1.12\%$ shift in consumption. Therefore, if anything, we find evidence for slightly more insurance to permanent shocks upon inclusion of the fourth-order moments. The dispersion and skewness of taste heterogeneity and consumption error do not change much from the lower-order cases; we find evidence for excess kurtosis in both cases, albeit at a much lesser extent than for income shocks. [^33] Appendix table C.7 uses imputed consumption from the CEX to target higher-order moments as in columns 3-4 of table 4. We do see slightly higher pass-through of permanent shocks upon inclusion of the higher-order moments, but we also find a strong, significant, and negative response to transitory shocks, which is hard to explain. We do not think those results are very credible; the imputation brings in substantial measurement/imputation error, which becomes worse the higher the order of the target moments is.
@@
Clearly, targeting higher-order moments in the context of the linear consumption function has not altered much the degree of partial insurance vis-à-vis targeting second-order moments alone. This is counterintuitive given the extent of tail income risk in the data, as well as the ample evidence in the literature pointing to the contrary. For example, Ai and Bhandari (2021) show that tail risk is uninsured in realistically parameterized labor markets, while Busch and Ludwig (2021) find sizable welfare implications of higher-order risk. [^34] De Nardi et al. (2019) and Busch and Ludwig (2021) find higher insurance when shocks are non-normal, due to stronger precautionary motives. The latter caution against BPP’s linear framework, showing evidence of bias in the presence of tail risk. Our quadratic specification addresses their concern. One would thus expect that tail risk increases the pass-through of shocks, as there must be a limit beyond which households cannot further insure consumption against extreme events.
Yet, the linear consumption function leaves no room to such events. BPP rely on a log-linearization of the consumption rule around a deterministic lifecycle profile (appendix A). But tail events move the agent far from that profile: if utility is sufficiently concave, a large shock will command a large consumption response. A linearized marginal utility, however, fails to pick up such large response (echoing Carroll, 2001), which in turn explains why we do not see much action upon including higher-order moments. In other words, tail risk has, by design, no role in the linear consumption function, which in turn motivates and justifies our higher-order specification subsequently.
@@
To summarize our findings from the linear consumption function: 1.) we replicate BPP over the recent years and find much lower pass-through of permanent shocks (larger degree of partial insurance); the imputation of consumption from the CEX and the biennial nature of the modern PSID explain the bulk of the difference from BPP; 2.) targeting higher-order moments does not alter the degree of consumption partial insurance, which is not entirely surprising as tail risk has, by design, no role in the linear framework.

### v4

@@ 4.2 Estimation
Estimation proceeds as follows. In a first step, we compute income and consumption growth net of their predictable components, $\mathbf{X}_{it}$ and $\mathbf{Z}_{it}$ in model notation, by regressing real income and consumption growth on a rich set of time effects and household characteristics. [^20] The controls include year and year-of-birth dummies, dummies for education, race, region, family size and number of children (level and change from last period), employment status (current and past), presence of outside dependents, and presence of income recipients other than the husband and wife (both current and past). The dummies for education, race, and region are interacted with the year dummies. The residuals from these regressions are the empirical counterparts to unexplained income and consumption growth, $\Delta y_{it}$ and $\Delta c_{it}$, respectively. In a second step, we estimate the parameters of the income process – i.e., the second and higher moments of permanent and transitory shocks – using income data alone. In a third step, we estimate the transmission parameters of shocks and the moments of taste heterogeneity and consumption measurement error.
@@
Unless noted otherwise, we conduct inference for the main parameters using block bootstrap. We thus account for serial correlation and heteroskedasticity of arbitrary form and for parameter uncertainty from earlier steps of the estimation process. GMM (asymptotic) standard errors clustered at the household level are very similar, suggesting that earlier-step parameter uncertainty is negligible (the income parameters are tightly estimated). [^22] We compute bootstrap standard errors as the normalized interquartile range of bootstrap replications, i.e., the difference between the $75^{\text{th}}$ and $25^{\text{th}}$ percentiles divided by the interquartile range of the standard normal cdf. This is equivalent to applying a normal approximation to the distribution of bootstrap replications, thereby shielding the standard errors from extreme resampling draws (Blundell et al., 2016; Theloudis, 2021). Bootstrap standard errors without this approximation are similar. For the subsample analysis (tables 5-6 and corresponding tables in the appendix), we calculate GMM (asymptotic) standard errors clustered at the household level to avoid numerical instability in bootstrap, driven by vanishing shares of certain subsamples, e.g., the wealthy non-college graduates, in many bootstrap draws. For the replication of BPP using the original BPP data (table C.5), we show non-clustered GMM standard errors, as in BPP.
@@ 4.3 Empirical moments of income and consumption
Table 1 reports the empirical second, third, and fourth moments of the cross-sectional distributions of unexplained income and consumption growth in the sample. [^24] In contrast to the identifying statements of section 3, the data are biennial so we report moments for income/consumption growth at a biennial frequency. We maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that this corresponds to a difference over two years. See appendix B for details.
Income moments. Panel A presents the moments of income growth. The variance is $0.153$, while the first-order autocovariance is $-0.049$ and highly significant. The income process of section 2.1 implies that income growth is negatively first-order autocorrelated due to mean reversion in the transitory shock, and this is supported by the data. [^25] The second-order autocovariance is an order of magnitude smaller and only marginally significant. This feature and the biennial nature of the data, whereby the second-order autocovariance reflects correlations 4 calendar years apart, motivates our white noise specification for the transitory shock.
@@
Notes: The table presents the second, third, and fourth moments of income and consumption growth. Skewness and kurtosis correspond to the third and fourth standardized moments respectively. Column 1 reports moments of consumption internally in the PSID; column 2 reports moments of consumption imputed from the CEX. See appendix C.4 for details on the imputation. We maintain the notation $\Delta x_{t}$ for the first difference of variable $x$ over time, noting that, given the biennial nature of the data, this corresponds to a difference over two calendar years. Block bootstrap standard errors are in parentheses.
@@
Guvenen et al. (2021) focus on male earnings while our model and statistics for income growth above revolve around household disposable income. To facilitate the comparison, appendix C.2 reports moments of male earnings growth in our sample. The variance is similar between the PSID and the administrative data, at $0.271$ and $0.323$ respectively. Skewness is $-0.569$ (and highly statistically significant) versus $-1.039$ in the administrative data. While the nature of skewness is similar in both cases, the population data feature substantially more left skewness than the PSID; Guvenen et al. (2021) confirm this. Kurtosis, finally, is nearly identical between the two, at $13.711$ and $13.494$ respectively.
Left skewness and excess kurtosis are thus not just features of the growth rates in disposable household income (table 1), but they are also observed in alternative measures of income, such as male earnings (table C.2, panel A) and household earnings (table C.2, panel B). Together, left skewness and excess kurtosis suggest that income changes for most households are either close to their mean value of zero or far out in the left tail (large income reductions). These results corroborate (and, in fact, are close numerically with) findings by De Nardi et al. (2019); De Nardi et al. (2021) over an earlier period of the PSID.
Large income reductions do not just reflect unemployment, but a multitude of events. Issues of reverse causality aside, appendix C.3 documents that households with large income reductions also experience the loss of a second job, worsening health, large drops in hours worked (reflecting both demand and supply factors), and precarious housing. It is unclear how persistent those events are, but our subsequent decomposition of income to moments of permanent and transitory shocks will shed light on this.
Consumption moments. Panel B of table 1 presents the moments of consumption growth; column 1 specifically reports moments from our baseline sample in the PSID. The variance of consumption growth is similar to that of income growth ($0.126$ versus $0.153$), reflecting either a large pass-through of income shocks or large taste heterogeneity/measurement error. The first-order autocovariance is $-0.044$ and highly significant. In the context of our model, this autocovariance reflects the variance of consumption error $\sigma_{u^{c}_{t}}^{2}$ (see appendix B).
@@
Appendix figure C.1 plots the variance, skewness, and kurtosis of log consumption in the PSID and CEX. The patterns for the variance and kurtosis are similar in the two samples over time, though the variance in the PSID is lower and the kurtosis is slightly higher. Skewness of log consumption, however, is markedly different: it is strongly positive in the PSID but mostly negative in the CEX; the patterns also diverge over time.
This comparison is not informative about the moments of consumption growth, which is the relevant variable in our case. The CEX lacks the panel dimension needed for our analysis, so it cannot be used to obtain measures of consumption growth. This is why BPP imputed consumption from the CEX into the PSID: the CEX lacked the necessary panel dimension while the PSID lacked comprehensive consumption data before 1999. We repeat BPP’s imputation over the recent years, i.e., our sample period 1999–2019. The imputation involves the estimation of food demand in the CEX (food expenditure expressed as a function of prices and nondurable consumption). The demand equation is then inverted (nondurable consumption expressed as a function of prices and food) and used to impute nondurables in the PSID based on food expenditure and prices that are available in both surveys. In this way, we obtain an alternative, external measure of consumption in the PSID, and of consumption growth. Appendix C.4 provides details on the imputation.
The moments of imputed consumption growth are in column 2 of table 1. The variance at $0.323$ is two-and-a-half times higher than the variance in the true series (column 1). In fact, the variance of imputed consumption growth is more than double the variance of income growth. The higher-order moments differ markedly between the imputed and true series. Skewness in imputed consumption growth is $-0.177$ (versus $-0.039$ in the true series), while kurtosis is $17.109$, almost four times higher than kurtosis in the true series. The imputation thus seems to impart substantial measurement/imputation error to consumption.
Since the imputation is not strictly needed over our sample period, we use consumption internally available in the PSID as our baseline measure of consumption. The present analysis also suggests that we may not want to rely on higher-order moments of consumption, given the discrepancies in skewness of log consumption between the PSID and the CEX, which are independent of the imputation. As we established in section 3.3, an advantage of our approach is precisely its non-reliance on higher-order consumption moments.
@@ 5 Results
We first present baseline results imposing stationarity in the distributions of income shocks, taste heterogeneity, consumption error, and a constant pass-through of income shocks over the lifecycle. We subsequently relax stationarity/lifecycle-invariance in section 5.3.282828Stationarity is not an identifying assumption but helps with precision, especially of higher moments.
@@ 5.1 Income process
We present estimates of the income process in three specifications. First, we target only the second moments of income growth, which enables us to estimate the variance of income shocks. This is similar to what BPP do over the earlier period 1980–1992. Second, we target second and third moments of income growth, which enables us to estimate also the skewness of shocks. Finally, we target up to the fourth moments, which allows us to estimate all the parameters of the income process as presented in section 2.1. The results are in table 2, while the full set of moments targeted in each case is shown in appendix B.4.1.
@@
Table 2: Estimates of the income process
@@
Variance $\sigma_{\zeta}^{2}$ 0.027 (0.002) 0.027 (0.002) 0.030 (0.002)
Skewness $\gamma_{\zeta}$ central -0.004 (0.003) -0.004 (0.003)
standardized -0.947 (0.649) -0.831 (0.571)
Kurtosis $\kappa_{\zeta}$ central 0.039 (0.007)
standardized 44.393 (11.080)
@@
Variance $\sigma_{v}^{2}$ 0.033 (0.002) 0.033 (0.002) 0.031 (0.003)
Skewness $\gamma_{v}$ central -0.008 (0.003) -0.008 (0.003)
standardized -1.310 (0.510) -1.431 (0.566)
Kurtosis $\kappa_{v}$ central 0.048 (0.005)
standardized 49.226 (10.119)
@@
Targeting the third moments of income growth (in addition to the second moments) leaves the variances unchanged. We estimate the skewness (third standardized moment) at $-0.947$ and $-1.31$ for permanent and transitory shocks, respectively. Both shocks thus exhibit substantial negative skewness, i.e., their distribution has a longer left than right tail. This is consistent with Geweke and Keane (2000) who find left skewness of earnings shocks in the early PSID, and Theloudis (2021) who finds left skewness of wage shocks in the recent period. These results vary with the level of income, but we postpone this discussion for later. Assuming for now that good and bad shocks transmit similarly into consumption, the long left tail implies that bad shocks are more damaging than good ones because they are further away from the zero mean – they are more extreme. This is true for both permanent and transitory shocks. Skewness of income growth, however, is driven mostly by permanent shocks (fully so under stationarity). Therefore, if targeting the third moments of income growth subsequently changes the degree of consumption insurance in the household, this must reflect the implications of permanent rather than of transitory shocks.
Targeting all moments of income growth up to fourth-order leaves the third central moments unchanged and modifies the variances by little. We estimate the kurtosis (fourth standardized moment) at $44.393$ and $49.226$ respectively. In both cases, the distributions are highly leptokurtic: they are sharp-pointed around the mean, have relatively little mass in the shoulders and very fat tails. This is consistent with Busch and Ludwig (2024), who use the PSID until 2012 and estimate numerically similar kurtosis coefficients for both shocks.
Seen together, large left skewness and excess kurtosis imply that most households experience either negligible shocks (shocks close to the zero mean) or shocks that are quite bad. Ignoring these stark features of income shocks may lead to erroneous inference of the degree of consumption partial insurance, to which we now turn.
@@ 5.2 Linear consumption function
We present estimates of the linear consumption function, first, targeting only second moments of the joint income-consumption distribution and, subsequently, targeting second as well as higher-order moments. The full set of targeted moments is shown in appendix B.4.2.
@@
What explains the discrepancy in $\phi^{(1)}$ with BPP? As suggested earlier, there are a few important differences between our replication and the original exercise in BPP: new consumption data and a different time period. To help decipher their role, we impute consumption from the CEX into the modern PSID in the same way BPP did for the earlier period, and we repeat the estimation. The results are in column 2 of table 3. The pass-through of permanent shocks doubles to $\phi^{(1)}=0.288$, while the pass-through of transitory shocks remains statistically zero (but at a larger magnitude). Using imputed consumption data thus inflates the transmission of permanent shocks and lowers the degree of insurance. This is likely the byproduct of measurement/imputation error, evident in the variance $\sigma^{2}_{u_{c}}$ of consumption error that more than triples between the true and the imputed data. By contrast, the variance of taste heterogeneity $\sigma^{2}_{\xi}$ remains similar in both cases, and very close to BPP. The imputation, therefore, seems to bring in large amounts of measurement error, explaining about 1/3 of the difference between our lower $\phi^{(1)}$ and BPP’s higher estimate. Appendix C.4 discusses the details of the imputation and a mechanism through which it inflates $\phi^{(1)}$ in particular (the imputation involves the inversion of a demand system, so certain types of measurement error may inflate the imputed income-consumption covariance – evident also in table 1).
Table 3: Estimates of the consumption function
@@
[…]
@@
Introducing higher-order moments. The estimation of the linear consumption function targeting higher-order moments is our first attempt to uncover how the tails of income risk affect the pass-through of shocks into consumption. In other words, we want to measure the degree of partial insurance targeting moments beyond the workhorse income-consumption covariance in order to account for the non-Gaussian features of modern income dynamics. We introduce the higher-order moments incrementally. [^32] We use the corresponding pre-estimated income parameters from columns 2 and 3 of table 2 respectively.
@@
Column 3 of table 3 presents the results from targeting third- in addition to second-order moments of consumption and income. The transmission parameters remain largely unchanged upon inclusion of the third moments. We estimate $\phi^{(1)}=0.160$ (as opposed to $0.152$ with only second moments) and $\psi^{(1)}=-0.012$ (again indistinguishable from zero). We thus again find large partial insurance to permanent shocks and full insurance to transitory shocks. Skewness in taste heterogeneity and consumption error is small (much smaller than skewness in income shocks), while their variances remain unchanged from the specification with only second moments. Therefore, left skewness in consumption is driven by negatively skewed income shocks rather than by taste heterogeneity or measurement error.
@@
Column 4 of table 3 presents the results from targeting all income and consumption moments up to fourth-order. The earlier findings hold up. We estimate $\phi^{(1)}=0.112$ (versus $0.152$ and $0.160$ in the lower-order cases) and $\psi^{(1)}=0.007$ (indistinguishable from zero). A $10\%$ permanent change in income thus results in $1.12\%$ shift in consumption. Therefore, if anything, we find evidence for slightly more insurance to permanent shocks upon inclusion of the fourth moments. The dispersion and skewness of taste heterogeneity and consumption error do not change much from the lower-order cases; we find evidence for excess kurtosis in both cases, albeit at a much lesser extent than for income shocks. [^33] Appendix table C.8 uses imputed consumption from the CEX to target higher-order moments as in columns 3-4 of table 3. We do see slightly higher pass-through of permanent shocks upon inclusion of the higher-order moments, but we also find a strong, significant, and negative response to transitory shocks, which is hard to explain. We do not think those results are very credible; the imputation brings in substantial measurement/imputation error, which becomes worse the higher the order of the target moments is.
@@
Clearly, targeting higher-order moments in the context of the linear consumption function has not altered much the degree of partial insurance vis-à-vis targeting second-order moments alone. This is counterintuitive given the extent of tail income risk in the data, as well as the ample evidence in the literature pointing to the contrary. [^34] Ai and Bhandari (2021) show that tail risk is uninsured in realistically parameterized labor markets. However, De Nardi et al. (2019) and Busch and Ludwig (2024) find higher insurance when shocks are non-normal, due to stronger precautionary motives. The latter caution against BPP’s linear framework, showing evidence of bias in the presence of tail risk. Our quadratic specification addresses their concern. One would expect that tail risk increases the pass-through of shocks, as there must be a limit beyond which households cannot further insure consumption against extreme events. Yet, the linear consumption function leaves no room to such events. BPP rely on a log-linearization of the consumption rule around a deterministic lifecycle profile (appendix A). But tail events move the agent far from that profile: if utility is sufficiently concave, a large shock will command a large consumption response. A linearized marginal utility fails to pick up such large response (Carroll, 2001), which in turn explains why we do not see much action upon including higher-order moments. In other words, tail risk has, by design, no role in the linear consumption function, which in turn motivates our higher-order specification subsequently.
@@
To summarize our findings from the linear consumption function: 1.) we replicate BPP over the recent years and find much lower pass-through of permanent shocks (larger degree of partial insurance); the imputation of consumption from the CEX and the biennial nature of the modern PSID explain the bulk of the difference from BPP; 2.) targeting higher-order moments does not alter the overall degree of consumption partial insurance, which is not entirely surprising as tail risk has, by design, no role in the linear framework.